OpenStudy (anonymous):

If u and v are orthogonal, what is the measure of the angle between them? 0° 45° 90° 180°

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@amistre64

OpenStudy (reynad):

orthogonal means perpendicular. Do you know what to do now?

OpenStudy (anonymous):

so its 90 degrees

OpenStudy (reynad):

Yes

OpenStudy (nincompoop):



OpenStudy (anonymous):

that angle would be 0?

OpenStudy (nincompoop):

dot product would be zero

OpenStudy (nincompoop):

gonal and normal vectors http://www.ucl.ac.uk/~ucahmdl/LessonPlans/Lesson10.pdf

OpenStudy (anonymous):

so It is 0

OpenStudy (michele_laino):

hint: 2 vectors are orthogonal each other when, their scalar product is zero, namely: \[\Large u \cdot v = \left\| u \right\|\;\left\| v \right\|\;\cos {\theta _{uv}} = 0\]

OpenStudy (nincompoop):

well you need to understand why zero

OpenStudy (anonymous):

well now I know that is definitely 0

OpenStudy (michele_laino):

being \[\Large {\theta _{uv}}\] the angle between such vectors

OpenStudy (nincompoop):

what you know does not matter, you need to show why it is zero

OpenStudy (anonymous):

I found the answer by finding the scalar product at first